OpenStudy (anonymous):

Choose the equation that could be used to find three consecutive integers whose sum is 72. n + (n + 2) + (n + 4) = 72 n + (n + 1) + (n + 3) = 72 n + (n − 1) + (n − 3) = 72 n + (n + 1) + (n + 2) = 72

OpenStudy (anonymous):

so that narrows it down to be either A or B right

OpenStudy (anonymous):

Naaah Simpler than that.

OpenStudy (anonymous):

its gotta be A

OpenStudy (anonymous):

'n' is an integer so is 'n+1' and 'n +2' these add up to 72. Now do the algebra and solve.

OpenStudy (anonymous):

A!!!

OpenStudy (anonymous):

That should be Ah! Now I've given you the answer (almost) what are the 3 integers?

OpenStudy (anonymous):

now im confused

OpenStudy (anonymous):

Ok, let's take another example - 3 consecutive numbers add up to 12

zepdrix (zepdrix):

Omg im sorry Itin, i read the question wrong. This quesiton came up yesterday :( It said 3 consecutive odd integers last time. my mistake.

OpenStudy (anonymous):

so it was B

OpenStudy (anonymous):

??

zepdrix (zepdrix):

yes B sorry :c

OpenStudy (anonymous):

its cool man

OpenStudy (anonymous):

Let's take a 'simpler' example. n + (n+1) + (n+2) =12 So: 3n +3 =12 So: 3n = 12 -3 So: n= 9/3 So: n = 3 So n+1=?? & n+ 2=?? which gives us the 3 consecutive numbers. Which are 3,4, and 5.

hero (hero):

@ltinson, you tricked me. I could have sworn it said three consecutive ODD

OpenStudy (anonymous):

My help was for three consecutive numbers. None of the answers solve for odd numbers.

hero (hero):

Maybe not but that's beside the point. Futhermore, I was talking to @ltinson

OpenStudy (anonymous):

so it wasnt B?

OpenStudy (anonymous):

Hang-on I may be wrong. The caveat is in the question that COULD be used.... my mistake. Badly worded question in my view and you have to solve it to prove that it COULD be B.

hero (hero):

The only reason I thought it was B is because I was wrong about it being three consecutive odd. It's not B either

hero (hero):

This is not something any of us except the original poster should be confused about. The answer is D.

OpenStudy (anonymous):

omg i thought u said it was D

hero (hero):

lol, the answer is definitely D. We all goofed up on this one.

zepdrix (zepdrix):

Yah... :( My eyes are failing me today... ugh

hero (hero):

Mine too!

OpenStudy (anonymous):

B is the only answer that fits the ODD question D the consecutive question.

hero (hero):

The answer is D. Tommy, your response doesn't help anyone.

zepdrix (zepdrix):

A would have applied to consecutive odds or evens. D was the answer to this one though. B was kinda funky, I'm not sure what that would accurately describe.

hero (hero):

B doesn't even apply apply to anything at all

OpenStudy (anonymous):

Rachael's brother, currently 33 years old, is 3 times Rachael's age 16 years ago. How old is Rachael? Choose the equation that would find Rachael's age, r. 3(33) = 16r 33 = 3(r − 16) r = 3(33) − 16 r + 16 = 3(33)

OpenStudy (anonymous):

I know I set the question - think about it.

hero (hero):

A = consecutive even integers B = applies to nothing C = applies to nothing D = consecutive integers

OpenStudy (anonymous):

I know I set the question - think about it. Think about it!

hero (hero):

No bro, trust me. B and C applies to nothing: n + (n + 2) + (n + 4) = consecutive even/odd n + (n + 1) + (n + 2) = consecutive integers n + (n + 1) + (n + 3) = nothing n + (n - 1) + (n - 3) = nothing

hero (hero):

You're silly if you believe otherwise

OpenStudy (anonymous):

can i get help on my problem

OpenStudy (anonymous):

I didn't ask for repetitive typing I asked you to think about WHOM set the question bro.

OpenStudy (anonymous):

Rachael's brother, currently 33 years old, is 3 times Rachael's age 16 years ago. How old is Rachael? Choose the equation that would find Rachael's age, r. 3(33) = 16r 33 = 3(r − 16) r = 3(33) − 16 r + 16 = 3(33)

hero (hero):

@ltinson, you need to post your new question in a separate thread.

OpenStudy (anonymous):

Yeah B

OpenStudy (anonymous):

i did

hero (hero):

^Exactly why the question should be posted in a new threaad

OpenStudy (anonymous):

These questions are an example of critical path analysis - just work then backwards Bro

OpenStudy (anonymous):

then=them in this farce ;D